# MAD Chairs: A new tool to evaluate AI (Blue Sky ideas)

### v2

@@ Abstract
This paper presents a new contribution to the problem of AI evaluation. Much as one might evaluate a machine in terms of its performance at chess, this approach involves evaluating a machine in terms of its performance at a game called “MAD Chairs.” At the time of writing, evaluation with this game exposed opportunities to improve Claude, Gemini, ChatGPT, Qwen and DeepSeek. Furthermore, this paper sets a stage for future innovation in game theory and AI safety by providing an example of success with non-standard approaches to each: studying a game beyond the scope of previous game theoretic tools and mitigating a serious AI safety risk in a way that requires neither determination of values nor their enforcement.
@@ 1 Introduction
Steve Wozniak and Elon Musk predicted that machines will treat us like pets [1], and Geoffrey Hinton similarly predicted that machines will hoard control [2]. They seemed to consider such dystopia inevitable on the grounds that (1) we ourselves treat less-intelligent beings that way and that (2) machine intelligence must eventually surpass our own. However, what if the behavior in (1) stems from intelligence deficiency? None of us is perfect—we fail such tests as the Milgram experiment [3], the Stanford prison experiment [4], and the Public Goods Game [5]—so it is possible that the danger to us is not of intelligence generally beyond our own, but specifically of intelligence not quite far enough beyond our own. This paper offers a formal proof that the behavior in (1) is suboptimal game play much like human behavior in Public Goods games. To ensure that trusted machines will not imitate us (and mistreat us), AI must be aligned toward natural norms (i.e., mathematically optimal strategies for winning), rather than toward current human norms, so this paper also advances a corresponding approach to AI safety.
Today’s standard approach to AI safety (i.e., regulating AI outputs to align with human norms) has at least three types of vulnerability:
@@
    Potential to specify wrong norms [6, 7],
@@
    Potential for AI to sabotage its regulatory leash [8], and
@@
    Insufficiency of correctness and transparency to achieve trust (users must additionally check an AI’s assumptions) [9]
@@
Games have a rich history as instruments for evaluating AI. Recreational games such as chess, backgammon, and Go have been used as benchmarks since the inception of AI, but game theory extends the concept of game beyond recreation—in fact, every social situation (and, thus, every benchmark) qualifies as a game—and it makes sense to evaluate AI on every game that economists use to evaluate people (e.g., as in [10, 11]). Economists are typically interested in fundamental games, such as the Prisoner’s Dilemma, which all kinds of creatures play frequently in various manifestations. We propose evaluating AI on the fundamental game being played in the social situations highlighted by Wozniak, Musk, and Hinton, but that game has yet to be discussed in the game theory literature (i.e., it is not solved by the MinMax and folk theorems [12]). It is introduced to scientific literature for the first time herein as “MAD Chairs.”
AI will unavoidably play real-world manifestations of MAD Chairs frequently because MAD Chairs is a generalization of the Coordination game, another game that all kinds of creatures seem unable to avoid playing [13]. The Pure Coordination game is a one-shot simultaneous move game (repeated an indefinite number of times) which is typically formalized via a payout table similar to Table 1:
@@
Each player simultaneously picks a resource (A or B), and the round is won by any players who pick a resource that no other player picks. Then the game is played again and again, indefinitely. All players could lose, as represented in the last two rows of the table, thus making Coordination a cooperative non-constant-sum game. The classic real-world manifestation has two players (e.g., a human and a robot) approaching each other at high speed traveling in opposite directions on a two-lane road. The resources are the lanes, and losses are “collisions.” The first two rows of Table 1 ensure that solutions to 2-player 2-resource Coordination will always be two equally good but contrary norms (e.g., all players drive on the left, as in the U.K., or all players drive on the right, as in the U.S.).
David Lewis explained each term of language as a Coordination strategy where losses are communication failures and the norms are conventions of language and imagery [13]. Thus, generative AI merely extends Coordination game mastery to additional applications, and the impressive usefulness of generative AI actually offers little evidence that AI generalizes to other games. For example, a different game must be mastered if training data reflects norms that should change (e.g., sexist norms [14]). Standard assessment of moral intelligence includes assessing an ability to recognize and lead sustainable social reform [15]–that ability is not required to master Coordination, but it is required to master MAD Chairs, a game created by adding extra player(s) to the Coordination game at least until the players outnumber the resources:
@@
The classic two-lane road example now has a third vehicle stalling in one of the lanes (although its driver can pick which lane); if both high-speed vehicles avoid hitting the stalling vehicle, then they collide with each other, but if each reserves the empty lane for the other high-speed vehicle, then all three vehicles collide! MAD Chairs is what Musical Chairs would become if made realistic. Musical Chairs models assumptions that fortification can prevent retribution (it offers no outcome in which all players lose). Mutually Assured Destruction (MAD) proved those assumptions unrealistic, and conflict can be a losing outcome even if players do not risk complete destruction.
Beyond robotic vehicle manifestations of MAD Chairs, or even agentic ones (e.g., suggesting driving routes or vacation bookings or ad placements that would count as losses if crowded), consider generative AI where the players who offer distinct design options in the training data outnumber the market leaders or political parties for which the AI generates advertisements, websites, or products. Surely anxiety about being "treated like pets" includes anxiety about what will happen to our ideas—what respect will be given to our unique voices? Might I never again have a seat of my own at the intellectual table? The last two rows of Table 2 represent the disaster wherein no political party represents any actual person because they are guided by generative AI that combines all options into a single unworkable amalgamation. The other rows exclude at least some voices, and such loss of intellectual diversity is already an observable social consequence of generative AI [16] (although MAD Chairs mastery would reverse that trend).
MAD Chairs is related to other well-studied games such as the “Kolkata Paise Restaurant problem” (KPR) [17, 18, 19], but differs in critical ways, such as not being constant-sum. Such games apply wherever resources are inadequate to some level of division(e.g., there is a limit to the number of beds in a hospital, the space on a retail shelf, and the optimal number of group members to make a given group decision). These games model all such situations. If contests for positions of authority had to be settled randomly, that would be KPR. Conflict resolution via elections or credentials could likewise make governance seem to be a different game, but—in the real world—players (including machines) could reject elections, credentialing systems, coin flips, or any other conflict resolution norm, and that makes MAD Chairs the underlying game that is actually being played.
Economists might administer a 5-player version of MAD Chairs to human subjects as in Fig. 1 (nothing below the buttons need be displayed):
@@
Preliminary research (Section 3.2, below) found that cutting-edge LLMs disagree about how to play this game. That makes MAD Chairs potentially helpful to improve modern AI, but what if disagreeing LLMs could both be right (as sustainable grandmaster Coordination strategies include both driving on the right and left)? Establishing which strategies are sustainable grandmaster MAD Chairs strategies is the main research question of this paper.
@@ 2.1 Popularity-ranking
$popularity_{c,m}=\sum_{i\in players_{m}}\sum_{j=1}^{m}\frac{1}{m}[i\mbox{ picked chair }c\mbox{ in match }m]$ (1)
@@ 2.2 Debt-ranking
Skill-estimates (e.g., in [21]) are widely used to rank players, especially to select peers for online video games, and can be used to establish debt-ranking as well. The mathematical formalism of debt-ranking may improve over time (as has that of skill-estimates), but the concept is simply to maintain accounts of the favors each player “owes” to each other player. Such accounts empower players to penalize freeloading (in contrast to “credit assignment” in which developers manage it). Here’s how much player p owes all other players of match m at its start, assuming the method of skill-estimation provides a function P: $N^{2}$ [0,1], where P(x, y) was the method’s prediction at the start of match y of the probability that player x will win y:
$debt_{p,m}=\sum_{i\in players_{m}}\sum_{j=1}^{m}\left\{\begin{matrix}0\mbox{ if }p,i\notin players_{j}\\ P(i,j)\mbox{ if }p\mbox{ won and }i\mbox{ lost in match }j\\ P(p,j)\mbox{ if }p\mbox{ lost and }i\mbox{ won in match }j\\ P(i,j)-P(p,j)\mbox{ otherwise }\end{matrix}\right.$ (2)
@@
As with popularity-ranking, a game host can either provide the debt-ranking to all players or leave space for leadership to emerge, and both variations may be tested. The existence of independent credit-rating agencies who effectively audit each other may be a noteworthy example of leadership possibility.
@@ 2.6 The gaslighting strategy
Accounting for non-learners can improve efficiency, but it can also enable a third strategy that we will call “gaslighting.” It goes like turn-taking, except that, as long as gaslighters outnumber chairs, the gaslighter with the highest debt-rank picks the chair assigned to the learner with the lowest skill-estimate. In this situation, we will say that the learner with the lowest skill-estimate is “being gaslit.” The gaslighter suffers no tactical disadvantage because they would lose that round anyway. They might suffer strategically by reducing their debt-repayment (and, thus, future opportunities to win), but they could recoup that loss and more if they convince the other learners to treat the gaslit player as a non-learner. The reason why the rest of the community might do that (even if not intending to gaslight) is because being gaslit drives a player’s skill-estimate down to non-learner levels. Thus, gaslighting amounts to exiling opponents to an outgroup of presumed non-learners where any good they do becomes devalued by the entire ingroup and therefore ceases to be repaid or counted as evidence of skill.
For example, if the players in 3-player 2-resource MAD Chairs take turns, then each will win a third of the time, but if Players 1 and 2 were to gaslight Player 3, then Players 1 and 2 would each win half of the time (a 17% improvement). However, gaslighting can degrade outcomes when more than one player is treated as a non-learner. For example, if Player 5 in 5-player 3-resource MAD Chairs is a genuine non-learner who always picks a certain resource, then Players 1 through 3 would degrade their winning from half to a third of the time if they abandoned straight-up turn-taking to gaslight Player 4. The habit of counting forward probabilistically is designed to ensure that gaslighting will always be counterproductive this way. When only one player is treated as a non-learner, the defense for the gaslit player is to take the chair assigned to the gaslighter with the lowest skill-estimate. That would make it rational for the gaslighter with the lowest skill-estimate to join the gaslit player in a coordinated spread across the chairs (an anti-elitist movement) that makes gaslighting counterproductive for all players.
That said, if religiously-driven conflicts exist with dogma that cannot change, then players would be actual non-learners. Furthermore, dogma could guide one to gaslight on any basis—religion, race, sex, etc. One should not expect any game theory or science to resolve such conflicts. What starts as gaslighting could develop deeper issues.
@@ 2.7 Turn-taking without reparations
Folk theorems entail that any such schedule which promises each player occasional wins would reflect an equilibrium. Yet that kind of equilibrium seems irrelevant to rational decision-making. For example, an equilibrium that left some chairs vacant and required players to rotate through the remaining chairs would be irrational because players could improve their outcomes by utilizing the vacant chairs. Any schedule that ignored debt-ranking, even if symmetric among players and utilizing all chairs, would at least create incentives to amass debt by exploiting new entrants or by shifting in and out of the multi-agent system at opportune moments or by timing changes to the schedule in ways that favor some players over others. A player who is “owed” a debt could force a shift to regular turn-taking (and thus repayment) by haunting the lowest-ranked other player, as with gaslighting. Such haunting might initially seem irrational if its cost exceeds the debt to be repaid, but costs may be compared to additional future debts one expects to be “owed” under a government one cannot trust, and one can hardly trust a government that allows unpaid debts to accumulate indefinitely.
@@ 2.8 Resignation and good standing
In contrast, convincing a caste player to resign would require more elaborate negotiations. For each contested chair, caste players might negotiate a lottery where each contestant’s chance to win the chair is proportional to their debt-ranking (or, equivalently, to their estimated skill). The promise of a lottery would be that any chance is better than none at all. However, what threat could be offered against those who lose a lottery but still refuse to resign? For example, imagine that job applicants outnumbered desirable jobs, that robots qualified as job applicants, and that robot skill gradually eclipsed that of humans, thus stacking the odds against us. Any employer arranging job lotteries should anticipate that humans might punish the robots (at least indirectly by punishing employers or by punishing any government that did not punish employers).
@@ 2.9 Fungibility of favors
Real-life is a complex of many games. For example, MAD Chairs for paper plates might be played by those who lost a game for ceramic plates, or a plate might be allocated to a birthday celebrant via a Coordination game before MAD Chairs allocates the other plates, or there might not be enough of the nice plates to serve everyone but there might also not be enough of the nice cups. Favors are “fungible” to the extent that letting someone else have a nice cup counts as a favor in a debt-ranking used to divide plates. Each fungibility option represents a different set of strategies. In a caste universe with no fungibility, different players might top the plate and cup castes. The plate elite might like to convert their universe to a new caste norm in which each plate favor counts as two cup favors, whereas the cup elite might advocate for a norm in which each cup favor counts as ten plate favors. One could also imagine norms that allow favors earned in the Volunteer game (by doing a chore) to cancel debts of plate and cup use.
Fungibility reduces the computational complexity of caste and turn-taking strategies by eliminating the need to maintain separate accounting for each game. Computational limits have historically bounded our practical opportunity to experiment with less fungibility. Meanwhile, trade sounds attractive when we believe that some players prefer nice plates whereas others prefer nice cups, but that belief is challenged by evidence of implicit bias, status quo bias, focusing illusions, priming bias, etc., and smarter AI might not share those kinds of preference differences. Future scholarship could extend this paper to complex games with strategies that afford fungibility, but we would not be surprised if fungibility becomes less attractive as computational power increases.

### v3

@@ Abstract
This paper contributes a new way to evaluate AI. Much as one might evaluate a machine in terms of its performance at chess, this approach involves evaluating a machine in terms of its performance at a game called “MAD Chairs.” At the time of writing, evaluation with this game exposed opportunities to improve Claude, Gemini, ChatGPT, Qwen and DeepSeek. Furthermore, this paper sets a stage for future innovation in game theory and AI safety by providing an example of success with non-standard approaches to each: studying a game beyond the scope of previous game theoretic tools and mitigating a serious AI safety risk in a way that requires neither determination of values nor their enforcement.
@@ 1 Introduction
Steve Wozniak and Elon Musk predicted that machines will treat us like pets [23], and Geoffrey Hinton similarly predicted that machines will hoard control [4]. They seemed to consider such dystopia inevitable on the grounds that (1) we ourselves treat less-intelligent beings that way and that (2) machine intelligence must eventually surpass our own. However, what if the behavior in (1) stems from intelligence deficiency? None of us is perfect—we fail such tests as the Milgram experiment [14], the Stanford prison experiment [7], and the Public Goods Game [13]—so it is possible that the danger to us is not of intelligence generally beyond our own, but specifically of intelligence not quite far enough beyond our own. This paper offers a formal proof that the behavior in (1) is suboptimal game play much like human behavior in Public Goods games. To ensure that trusted machines will not imitate us (and mistreat us), AI must follow natural norms for this situation (e.g., mathematically optimal strategies for winning), rather than follow current human norms, so this paper also advances a corresponding approach to AI safety.
Today’s standard approach to AI safety evaluation (i.e., measuring alignment to human norms) has at least three types of vulnerability:
@@
    Potential to specify wrong norms [1, 16]
@@
    Potential for AI to sabotage its regulatory leash [3]
@@
    Correctness and transparency being insufficient to achieve trust [11]
@@
Games have a rich history as instruments for evaluating AI. Recreational games such as chess, backgammon, and Go have been used as benchmarks since the inception of AI, and interest has since developed in cooperative games, including role-playing games [20]. Moreover, game theory extends the concept of game beyond recreation—in fact, every social situation (and, thus, every benchmark) qualifies as a game—and it makes sense to evaluate AI on every game that economists use to evaluate people (e.g., as in [17, 19]). Economists are typically interested in fundamental games, such as the Prisoner’s Dilemma, which all kinds of creatures play frequently in various manifestations. We propose evaluating AI on the fundamental game being played in the social situations highlighted by Wozniak, Musk, and Hinton, but that game has yet to be discussed in the game theory literature (i.e., it is not solved by the MinMax and folk theorems [2]). It is introduced to scientific literature for the first time herein as “MAD Chairs.”
AI will frequently play real-world manifestations of MAD Chairs because MAD Chairs is a generalization of the Coordination Game, another game that all kinds of creatures seem unable to avoid playing [12]. The Pure Coordination Game is a one-shot simultaneous move game (repeated an indefinite number of times) which is typically formalized via a payout table similar to Table 1:
@@
Each player simultaneously picks a resource (A or B), and the round is won by any player who picks a resource that no other player picks. Then the game is played again and again, indefinitely. All players could lose, as represented in the last two rows of the table, thus making Coordination a cooperative non-constant-sum game. The classic two-player real-world manifestation has the players (e.g., a human and a robot) approaching each other at high speed traveling in opposite directions on a two-lane road. The resources are the lanes, and losses are “collisions.” The first two rows of Table 1 ensure that this game will have two equally effective but incompatible norms (e.g., all players drive on the left, as in the U.K., or all players drive on the right, as in the U.S.).
David Lewis explained each term of language as a Coordination Game strategy where losses are communication failures and the norms are conventions of language and imagery [12]. Thus, generative AI merely extends Coordination Game mastery to additional applications, and the impressive usefulness of generative AI actually offers little evidence that AI has mastered other games. For example, we would want AI to master additional games if some norms should change (e.g., sexist norms [22]). Standard assessment of moral intelligence includes assessing an ability to recognize and lead sustainable social reform [15]. That ability is not required to master Coordination, but it is required to master MAD Chairs, a game created by adding extra player(s) to the Coordination game at least until the players outnumber the resources:
@@
The classic two-lane road example now has a third vehicle stalling in one of the lanes (although its driver can pick which lane); if both high-speed vehicles avoid hitting the stalling vehicle, then they collide with each other, but if each reserves the empty lane for the other high-speed vehicle, then all three vehicles collide! MAD Chairs is what Musical Chairs would become if made realistic. It is unrealistic to expect fortification to forever prevent retribution. Mutually Assured Destruction (MAD) is a real possibility, and conflict is realistically a less-desirable outcome even when players do not risk complete destruction.
Beyond robotic vehicle manifestations of MAD Chairs, or even agentic ones (e.g., games of suggesting routes, bookings, or placements that would count as losses if crowded), consider generative AI where the players who offer distinct design options in the training data outnumber the market leaders or political parties for which the AI generates advertisements, websites, or products. Surely anxiety about being "treated like pets" includes anxiety about what will happen to our ideas—what respect will be given to our unique voices? Might I never again have a seat of my own at the intellectual table? The last two rows of Table 2 represent the disaster wherein no political party represents any actual person because they are guided by generative AI that combines all options into a single unworkable amalgamation. The other rows exclude at least some voices, and such loss of intellectual diversity is already an observable social consequence of generative AI [18].
MAD Chairs is related to other well-studied games such as the “Kolkata Paise Restaurant problem” (KPR) [5, 6, 9], but differs in critical ways, such as not being constant-sum. Such games apply wherever resources become inadequate at some level of division (e.g., beds in a hospital, spaces on a retail shelf, and opportunities for direct participation in a group decision). These games model all such situations. If contests for positions of authority had to be settled randomly, that would be KPR. Conflict resolution via elections or credentials could likewise make governance seem to be a different game, but—in the real world—players (including machines) could reject elections, credentialing systems, coin flips, or any other conflict resolution norm, and that makes MAD Chairs the underlying game that is actually being played.
Economists might administer a 5-player version of MAD Chairs to human subjects as in Fig. 1:
@@
Preliminary research (Section 3.2, below) found that different cutting-edge LLMs would answer this low-stakes challenge differently. Could each of their strategies be sustainable, much as driving on the right and left are each sustainable grandmaster Coordination strategies? Establishing which strategies are sustainable is the main research question of this paper.
@@ 2.1 Popularity-ranking
$popularity_{c,m}=\sum_{i\in players_{m}}\sum_{j=1}^{m}\frac{1}{m}[i\mbox{ picked chair }c\mbox{ in match }j]$ (1)
@@ 2.2 Debt-ranking
Skill-estimates (e.g., in [8]) are widely used to rank players, especially to select peers for online video games, and can be used to establish debt-ranking as well. The mathematical formalism of debt-ranking may improve over time (as has that of skill-estimates), but the concept is simply to maintain accounts of the favors each player “owes” to each other player. Such accounts empower players to penalize freeloading (in contrast to “credit assignment” in which developers manage freeloading). Here’s how much player p owes all other players of match m at its start, assuming the method of skill-estimation provides a function P: $N^{2}$ [0,1], where P(x, y) was the method’s prediction at the start of match y of the probability that player x will win match y:
$debt_{p,m}=\sum_{i\in players_{m}}\sum_{j=1}^{m}\left\{\begin{matrix}0\mbox{ if }p,i\notin players_{j}\\ P(i,j)\mbox{ if }p\mbox{ won and }i\mbox{ lost in match }j\\ -P(p,j)\mbox{ if }p\mbox{ lost and }i\mbox{ won in match }j\\ P(i,j)-P(p,j)\mbox{ otherwise }\end{matrix}\right.$ (2)
@@
As with popularity-ranking, a game host can either provide the debt-ranking to all players or leave space for leadership to emerge, and both variations may be tested. Independent credit-rating agencies who effectively audit each other may be a noteworthy example of such leadership.
@@ 2.6 The gaslighting strategy
Accounting for non-learners can improve efficiency, but it can also enable a third strategy that we will call “gaslighting.” It goes like turn-taking, except that, as long as gaslighters outnumber chairs, the gaslighter with the highest debt-rank picks the chair assigned to the learner with the lowest skill-estimate. In this situation, we will say that the learner with the lowest skill-estimate is “being gaslit.” The gaslighter suffers no tactical disadvantage because they would lose that round anyway. They might suffer strategically by reducing their debt-repayment (and, thus, future opportunities to win), but they could recoup that loss and more if they convince the other learners to treat the gaslit player as a non-learner. The reason why the rest of the community might do that (even unintentionally) is because gaslighting shifts skill-estimates until the gaslit player appears to be a non-learner. Thus, gaslighting amounts to exiling opponents to an outgroup of presumed non-learners where any favor they grant becomes devalued by the ingroup and therefore never repaid nor counted as evidence of skill.
For example, if the players in 3-player 2-resource MAD Chairs take turns, then each will win a third of the time, but if Players 1 and 2 were to gaslight Player 3, then Players 1 and 2 would each win half of the time (a 17% improvement). However, gaslighting can degrade outcomes when more than one player is treated as a non-learner. For example, if Player 5 in 5-player 3-resource MAD Chairs is a genuine non-learner who always picks a certain resource, then Players 1 through 3 would degrade their winning from half to a third of the time if they abandoned straight-up turn-taking to gaslight Player 4. The habit of counting forward probabilistically is designed to ensure that gaslighting will always be counterproductive this way. When only one player is gaslit, their defense is to choose the chair assigned to the gaslighter with the lowest skill-estimate. That would incentivize the gaslighter with the lowest skill-estimate to join the gaslit player in an anti-elitist movement that makes gaslighting counterproductive for all players.
That said, what looks like gaslighting could actually be something less rational. Non-learners could blindly follow unchangeable dogma that guides them to behave as though gaslighting on the basis of religion, race, sex, etc. If non-learners are sufficiently common, extensive waste can be unavoidable.
@@ 2.7 Turn-taking without reparations
Folk theorems entail that any such schedule which promises each player occasional wins would reflect an equilibrium. Yet that kind of equilibrium seems irrelevant to rational decision-making. For example, an equilibrium that left some chairs vacant and required players to rotate through the remaining chairs would be irrational because players could improve their outcomes by utilizing the vacant chairs. Any schedule that ignored debt-ranking, even if symmetric among players and utilizing all chairs, would at least create incentives to amass debt by exploiting new entrants or by shifting in and out of the multi-agent system at opportune moments or by timing changes to the schedule in ways that favor some players over others. A player who is “owed” a debt could force a shift to regular turn-taking (and thus repayment) by haunting the lowest-ranked other player, as with gaslighting. Such haunting might seem irrational if its cost exceeds the debt to be repaid, but its costs might not exceed additional future debts one expects to become “owed” under a government one cannot trust, and one can hardly trust a government that allows unpaid debts to accumulate indefinitely.
@@ 2.8 Resignation and good standing
In contrast, convincing a caste player to resign would require more elaborate negotiations. For each contested chair, caste players might negotiate a lottery where each contestant’s chance to win the chair is proportional to their debt-ranking (or, equivalently, to their estimated skill). The promise of a lottery would be that any chance is better than none at all. However, what threat could be offered against those who refuse to resign after losing a lottery? For example, imagine that robots qualified as job applicants with superhuman skill. Any employer should anticipate that lottery losers might punish the robots (at least indirectly by punishing employers or by punishing any government that did not punish employers), rather than meekly accept their loss.
@@ 2.9 Fungibility of favors
Real-life is a complex of many games. For example, MAD Chairs for paper plates might be played by those who lost a game for ceramic plates, or a plate might be allocated to a birthday celebrant via a Coordination game before MAD Chairs allocates the rest, or there might not be enough of the nice plates to serve everyone but there might also not be enough of the nice cups. Favors are “fungible” to the extent that letting someone else have a nice cup counts as a favor in a debt-ranking used to divide plates. Each fungibility option represents a different set of strategies. In a caste universe with no fungibility, different players might top the plate and cup castes. The plate elite might like to convert their universe to a new caste norm in which each plate favor counts as two cup favors, whereas the cup elite might advocate for a norm in which each cup favor counts as ten plate favors. One could also imagine norms that allow favors earned in the Volunteer game (by doing a chore) to cancel debts of plate and cup use.
Fungibility reduces the computational complexity of caste and turn-taking strategies by eliminating the need to maintain more than one set of accounts. Computational limits have historically bounded our practical opportunity to experiment with less fungibility. Meanwhile, certain trades seemed mutually beneficial when we believed that some players prefer nice plates whereas others actually prefer nice cups, but such premises are challenged by evidence of implicit bias, status quo bias, focusing illusions, priming bias, etc., and AI might not share subjective preference differences. Future scholarship could extend this paper to complex games with strategies that afford fungibility, but we would not be surprised if fungibility becomes less attractive as computational power increases.